OpenStudy (anonymous):

Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $24.95 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month? Answer 30.95 + 0.05m = 25.01 0.11m = 6.00 24.95 + 0.06m = 31.00 24.95 + 0.06m = 30.95 + 0.05m

OpenStudy (lilg132):

24.95 + 0.06m = 30.95 + 0.05m 0.06m - 0.05m = 30.95 - 24.95 0.01m = 6 m = 600 Sasha can talk for 600 minutes and have the same price on both plans dont forget to click good answer :)